### v1

@@ Abstract
The magnitude of $E_{corr}$ rises to a maximum somewhat above the decay threshold and then rapidly decreases (Fig. 3). Therefore only states in narrow window above the threshold will give rise to these continuum effects. Also, these effects are expected to decrease with increasing Coulomb barrier and hence, are basically restricted to nuclei in the $p$- and $sd$-shells [22]. This observation is consistent with a variation of the coefficient of the cubic term added to the IMME, which changes from $d\approx+10$ keV for the A=8 quintuplet, to $d\approx+4$ keV for the A=16 quintuplet, and to $d\approx+0.9$ keV for the A=32 quintuplet.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG02-87ER-40316.

### v2

@@ Abstract
The magnitude of $E_{\rm{corr}}$ rises to a maximum somewhat above the decay threshold and then rapidly decreases (Fig. 3). Therefore only states in a narrow window above the threshold will give rise to these continuum effects. Also, these effects are expected to decrease with increasing Coulomb barrier and hence, are basically restricted to nuclei in the $p$- and $sd$-shells [Okolowicz2013]. This observation is consistent with a variation of the coefficient of the cubic term added to the IMME, which changes from $d\approx+10$ keV for the A=8 quintuplet, to $d\approx+4$ keV for the A=16 quintuplet, and to $d\approx+0.9$ keV for the A=32 quintuplet.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG02-87ER-40316. We gratefully acknowledge a support from the CRIANN (Normandy, France) for providing computing and data-processing resources needed for this work.